OpenStudy (anonymous):

how do i solve this x-y=4 and 2x-2y=4 please help me :(

OpenStudy (anonymous):

Solve one equation for one of the variables, plug that into the other equation to get a number value, then plug that number value back into the original equations to get the second variable.

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.

OpenStudy (anonymous):

Since -y does not contain the variable to solve for, move it to the right-hand side of the equation by adding y to both sides. x=y+4

OpenStudy (anonymous):

Since -2y does not contain the variable to solve for, move it to the right-hand side of the equation by adding 2y to both sides. 2x=2y+4 Divide each term in the equation by 2 2x/2 =2y/2 + 4/2 simplify x=y+2